Circular No. 7020

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1550–564 R. Jain and C. Bailyn, Yale University; and J. Orosz, Pennsylvania

State University, report YALO consortium observations using the Yale 1-m telescope at the Cerro Tololo Interamerican Observatory: “We have monitored the optical counterpart to XTE J1550–564 (IAUC 7009) nightly since Sept. 13 UT. The source has varied by < 0.3 mag in V since its discovery. We have adjusted our instrumental magnitudes so that V = 16.7 and V − I = +1.80 near Sept. 16.0, as reported by Castro-Tirado et al. (IAUC 7013), and we find that 16.6 < V < 16.9 over the range Sept. 8.99–21.97. Thus there is no dramatic optical response to the large increase in the x-ray intensity over the last several days (cf. IAUC 7019). Between Sept. 15.0 and 21.0, there was a trend in the V − I color where the source became increasingly redder. We find V − I = 1.80 ± 0.04 near Sept. 15.0 and 1.92 ± 0.01 near Sept. 21.0. YALO consortium monitoring observations of XTE J1550–564 will continue; additional observations are encouraged.”

COMET C/1998 M5 (LINEAR)

Ephemeris extension to IAUC 7004 (elements MPC 32410):

1998TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Sept. 24 19^{h}58.^{m}56 +40^{o}17.^{′}0 1.764 2.331 111.^{o}9 23.^{o}5 10.9
Oct. 4 19 33.49 +39 31.0 1.803 2.253 103.2 25.6 10.8
14 19 14.39 +38 27.3 1.855 2.178  94.7 27.2 10.7
24 19 00.91 +37 24.1 1.910 2.106  86.9 28.1 10.6
Nov. 3 18 52.20 +36 34.3 1.961 2.039  80.0 28.6 10.6
13 18 47.31 +36 05.7 2.002 1.976  74.2 28.8 10.5
23 18 45.45 +36 03.5 2.029 1.920  69.5 28.8 10.4
Dec. 3 18 45.95 +36 31.4 2.038 1.870  66.1 28.8 10.3
13 18 48.29 +37 32.9 2.028 1.827  64.1 29.0 10.2
23 18 52.10 +39 12.0 1.998 1.793  63.6 29.4 10.0

Visual m_{1} estimates: Aug. 26.85 UT, 10.8 (V. Znojil, Brno, Czech Republic, 25×100 binoculars); Sept. 1.00, 11.1 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); 8.85, 10.6 (M. Reszelski, Szamotuly, Poland, 0.25-m reflector); 15.92, 11.1 (J. Carvajal, Madrid, Spain, 0.32-m reflector); 19.04, 11.5 (J. Bortle, Stormville, NY, 0.41-m reflector).

1998 September 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green